Solve for X 1.42 = 1.55(1+ 4.5%)-X
Answer #1

1.42 = 1.55 * (1 + 4.5%) - X

1.42 = 1.55 * (1 + 0.045) - X

1.42 = 1.55 * (1.045) - X

1.42 = 1.61975 - X

X = 1.61975 - 1.42

X = 0.19975